The I.Sc lecture theatre of a college is 40 ft wide and has a door at a corner. A teacher enters at 12.00 noon through the door and makes 10 rounds along the 40 ft wall back and forth during the period and finally leaves the classroom at 12.50 p.m. through the same door. Compute his average speed and average velocity.
1
Expert's answer
2019-08-19T09:44:58-0400

Time consumed by teacher in room t = 12.50-12= 50 min = 3000 s

teacher covers in the room in one round= 40 + 40 = 80 ft

Thus for 10 rounds, the distance covered is

= 10 × 80 = 800 ft

Now, average speed = total distance/total time = 800/3000 = (4/15) ft/s

For average velocity, formulae deployed include displacement but not the distance and given as displacement/time.

Since the teacher comes and go with the same room, thus displacement is zero and thus the average velocity is zero